OpenStudy (anonymous):

What are the odds of rolling a 6 on any of 3 dice?

OpenStudy (anonymous):

1/2 ?

OpenStudy (geometry_hater):

3/18 I dont know depends on how many sides the dice has

OpenStudy (anonymous):

6-sided dice here, and we're looking for a six on any of the three, including if you get more than one 6

OpenStudy (anonymous):

Funny that a six on all of them is easy, right, but a six on any of them is a stumper.

OpenStudy (anonymous):

1 minus the odds of not throwing any 6.\[1-\left(\frac{5}{6}\right)^3 = \frac{91}{216} \]